OpenStudy (anonymous):

Factor completely: 49n2+14n+1?

OpenStudy (anonymous):

49n^2+7n +7n +1=0...can u do it nw

OpenStudy (anonymous):

\[49n^2+14n+1=(7n)^2+2\times 7n\times 1+1^2\]does this help???

OpenStudy (anonymous):

\[7n(7n+1) +1(7n+1) =0\] can u do it nw